OpenStudy (keana):

Find the vertex, axis of symmetry, domain, and range of the graph of y = -2x2 - 6x + 1.

OpenStudy (ash2326):

@keana have you studied Parabolas?

OpenStudy (keana):

eh... @ash2326

OpenStudy (anonymous):

you have to put the equation in vertex form: \[y=a(x-h)^2+k\]\[y=-2(x+\frac{3}{2})^2+\frac{11}{2}\]\[\] Vertex: (-\frac{3}{2}, \frac{11}{2}) domain is all real numbers

OpenStudy (anonymous):

i agree with @ash2326 this a parabolas problem

OpenStudy (anonymous):

for the axis of symmetry you use the -b/2a that come from the quadratic equation -b+- sqrt(b^2-4ac)/(2a)

OpenStudy (anonymous):

a=-2 don't need to uses to find the axis of symmetric b=-6 c=1

OpenStudy (anonymous):

Now uses the formula to find the axis of symmetry

OpenStudy (keana):

@purplec16 what does "\frac{11}{2})" mean

OpenStudy (anonymous):

It was unintentional it's latex rapping for when using the equation button in your post click equation put it in and see what happens :)

OpenStudy (anonymous):

Vertex:\[(-\frac{3}{2}, \frac{11}{2})\]